OpenStudy (anonymous):

Factor: 25m^8 - 60m^4n + 36n^2

OpenStudy (anonymous):

25m^8 - 60m^4n + 36n^2 This can be rewritten as (5m^4)^2 - 2(5m^4)(6n) + (6n)^2 This is of the form a^2 - 2ab + b^2 and hence using identity a^2 - 2ab + b^2 = (a - b)^2 (5m^4)^2 - 2(5m^4)(6n) + (6n)^2 = (5m^4 - 6n)^2 so 25m^8 - 60m^4n + 36n^2 = (5m^4 - 6n)^2